### v1

# Superficial Self-Improved Reasoners Benefit from Model Merging

Abstract
As scaled language models (LMs) approach human-level reasoning capabilities, self-improvement emerges as a solution to synthesizing high-quality data corpus. While previous research has identified model collapse as a risk in self-improvement, where model outputs become increasingly deterministic, we discover a more fundamental challenge: the superficial self-improved reasoners phenomenon. In particular, our analysis reveals that even when LMs show improved in-domain (ID) reasoning accuracy, they actually compromise their generalized reasoning capabilities on out-of-domain (OOD) tasks due to memorization rather than genuine learning. Through a systematic investigation of LM architecture, we discover that during self-improvement, LM weight updates are concentrated in less reasoning-critical layers, leading to superficial learning. To address this, we propose Iterative Model Merging (IMM), a method that strategically combines weights from original and self-improved models to preserve generalization while incorporating genuine reasoning improvements. Our approach effectively mitigates both LM collapse and superficial learning, moving towards more stable self-improving systems.
[…]
Superficial Self-Improved Reasoners Benefit from Model Merging
Xiangchi Yuan1, Chunhui Zhang2, Zheyuan Liu3, Dachuan Shi1, Soroush Vosoughi2, Wenke Lee1 1Georgia Institute of Technology, 2Dartmouth College, 3University of Notre Dame {xyuan300, dshi77, wenke}@gatech.edu {chunhui.zhang.gr, soroush}@dartmouth.edu zliu29@nd.edu
1 Introduction
The remarkable reasoning capabilities Jaech et al. (2024); Zhang et al. (2024); Guo et al. (2025) of large language models (LLMs) largely benefits from vast amounts of high-quality reasoning data. However, as the data corpus runs out Sutskever (2024) and increasingly powerful models approach human-level intelligence DeepMind (2024a, b), pressing issues emerge: (i) How to advance models’ reasoning capabilities despite data scarcity? (ii) How to obtain training data that exceeds human-level performance for next-generation models? A promising answer to both questions is model self-improvement or self-evolution, where models autonomously generate infinite high-quality data, which potentially surpasses human annotations, to continuously enhance their own performance.
Figure 1: The Superficial Self-Improved Reasoners phenomenon is mitigated by iterative model merging. Our method improves ID and OOD reasoning performances.
Although self-improvement has achieved remarkable success in specific domains such as mathematics OpenAI (2025); DeepMind (2024a), coding Li et al. (2022), and games Hu et al. (2024); Silver et al. (2018), recent studies reveal significant risks associated with using self-generated synthetic data for fine-tuning: in particular, model performance can degrade over multiple iterations of self-improvement, a phenomenon known as model collapse. Shumailov et al. (2023). In current research, model collapse is primarily attributed to a reduction in sampling diversity Shumailov et al. (2023); Alemohammad et al. (2024); Guo et al. (2024). To mitigate this problem, several studies suggest refreshing synthetic data with real data Bertrand et al. (2024); Alemohammad et al. (2024), accumulating data across training steps Gerstgrasser et al. (2024), and incorporating data verifiers Gillman et al. (2024) or correctors Feng et al. (2025). However, by focusing solely on data quality and diversity, these approaches overlook a more critical question: whether self-improvement genuinely enhances reasoning capabilities or merely memorizes the training distribution. This distinction becomes crucial when considering the model’s ability to generalize beyond its training data.
In this paper, we investigate a fundamental risk in model self-improvement for reasoning tasks that deepens the known challenge of model collapse. We identify a phenomenon we call Superficial Self-Improved Reasoners, where models appear to improve but actually fail to develop genuine reasoning capabilities. While these models show enhanced performance on in-domain (ID) reasoning tasks, they significantly underperform on out-of-domain (OOD) tasks, suggesting memorization rather than genuine reasoning improvement. To understand the mechanistic cause of this phenomenon, we perform a systematic analysis of the model architecture during self-improvement. By examining layer importance and parameter changes, we uncover a critical mismatch: the largest weight updates occur in layers that contribute least to reasoning, while reasoning-critical layers receive minimal updates. This mismatch explains why models tend to memorize training patterns rather than develop generalizable reasoning skills. To address this issue, we propose Iterative Model Merging (IMM), a novel method that strategically combines weights from original and self-improved models. IMM specifically targets the layer misalignment problem by preserving the stability of reasoning-critical layers while allowing beneficial updates from self-improvement. As demonstrated in Figure 1, this approach effectively balances performance improvements with preserved generalization ability.
A summary of the contributions is given below:
  • This work identifies the risk of self-improvement for reasoning: while the model enhances its reasoning capabilities, it still tends to memorize the training data, resulting in a loss of generalized reasoning ability. We refer to this phenomenon as Superficial Self-Improved Reasoners.
  • We provide an explanation for the emergence of Superficial Self-Improved Reasoners by highlighting a mismatch between the reasoning-critical layers and the layers that undergo the largest weight changes during self-improvement.
  • We propose IMM (Iterative Model Merging) to mitigate Superficial Self-Improved Reasoners. IMM offers an efficient approach to integrate the reasoning improvements of the self-improved model while preserving the generalization ability of the original model.
Figure 2: Superficial Self-improved Reasoners. The model’s performance is only improved on in-domain reasoning datasets while losing the generalized reasoning capabilities on out-of-domain reasoning datasets.
2 Related Work
LLMs for Reasoning
LLMs have demonstrated remarkable success across various reasoning tasks, including mathematical problem-solving, code generation, common-sense reasoning, and multimodal understanding Yu et al. (2024b); Lewkowycz et al. (2022); Wang et al. (2023); Zhang et al. (2025a, b). Beyond leveraging sophisticated prompting techniques to enhance reasoning capabilities Kojima et al. (2022); Wei et al. (2022); Zheng et al. (2023); Yao et al. (2024), many methods focus on fine-tuning LLMs with reasoning datasets to create more robust reasoners Lu et al. (2024); Yu et al. (2024b). For instance, approaches like SI Huang et al. (2022), STaR Zelikman et al. (2022), V-STaR Hosseini et al. (2024), and rSTaR Qi et al. (2025) fine-tune LLMs on task-specific datasets or synthesize reasoning data tailored for corresponding tasks. In addition to training models to generate correct answers, some studies introduce external verifiers Cobbe et al. (2021); Lightman et al. (2023); Hosseini et al. (2024); Yuan et al. (2024a) that select the best answer from a set of candidate solutions.
LLM Self-Improvement
Given the high cost of labeling data, it is increasingly common to leverage LLMs to generate synthetic responses for training student models. Traditionally, this process has focused on knowledge distillation from stronger teacher models Yuan et al. (2023); Wu et al. (2024). More recently, studies have demonstrated that distilling from weaker models—referred to as weak-to-strong knowledge distillation—can be more beneficial for LLMs compared to distilling from stronger models, given the same computational budget Bansal et al. (2024). Another emerging direction is LLM self-improvement, where models improve themselves using their own outputs Huang et al. (2022); Gulcehre et al. (2023); Singh et al. (2023). In the context of reasoning tasks, various self-improvement methods have been proposed: SPO Prasad et al. (2024) employs Self-Consistency Preference Optimization for self-improvement; Pang et al. (2024) iteratively generate and refine data to optimize the model’s reasoning ability; and Hosseini et al. (2024) utilize both correct and incorrect answers to improve reasoning performance through training an additional verifier.
Model Collapse
As real-world data becomes increasingly scarce Sutskever (2024), synthetic data is playing a crucial role in training modern generative models due to its low cost and infinite availability. However, recent studies have revealed the risks associated with this "free lunch," a phenomenon known as model collapse Shumailov et al. (2023). The model collapse has been extensively identified and analyzed in both computer vision Hataya et al. (2023); He et al. (2022); Bohacek and Farid (2023) and natural language processing Alemohammad et al. (2024); Liu et al. (2024); Gerstgrasser et al. (2024). Researchers have investigated its underlying causes from both empirical Padmakumar and He (2024); Dou et al. (2024); Guo et al. (2023) and theoretical perspectives Yuan et al. (2024a); Bertrand et al. (2023); Seddik et al. (2024). Current approaches to mitigating model collapse predominantly focus on data-centric methods. Feng et al. (2025) show that imperfect verifiers can help prevent model collapse by selecting appropriate data. Shumailov et al. (2023) proposes mixing data from previous iterations to prevent performance degradation, while Gerstgrasser et al. (2024) demonstrates that accumulating synthetic data over iterations reduces the risk of collapse.
3 Superficial Self-improved Reasoners
A natural and critical question arises for LLM self-improvement: does learning from synthetic reasoning data generated by the model itself trade off generalization ability for improved reasoning performance because of learning from itself? Our study shows that the answer is yes. In this section, we first confirm that self-improvement enhances in-domain reasoning performance but degrades general reasoning capabilities. We then investigate the underlying cause of this phenomenon by analyzing the layer-wise importance of the model during reasoning and tracking weight changes throughout the self-improvement process. A detailed comparison reveals a notable mismatch: the layers most crucial for reasoning experience relatively small weight updates, while less critical layers undergo more significant changes. This suggests that strong reasoning layers fail to substantially improve their reasoning ability through weight updates, whereas less important layers tend to overfit the training data rather than truly learning to reason.
3.1 Identify Superficial Self-improved Reasoners from OOD datasets
In this part, we identify Superficial Self-improved Reasoners by self-improving LLMs on the in-domain (ID) reasoning datasets and test them on out-of-domain (OOD) reasoning datasets.
Synthesizing Reasoning Data for Self-improvement
We begin by establishing the self-improvement framework through the generation of reasoning data. Following prior work Zelikman et al. (2022), we first synthesize reasoning data for fine-tuning. Let $\mathcal{D}=\{(q_{i},a_{i})\}_{i=1}^{n_{d}}$ denote a training dataset containing $n_{d}$ reasoning questions $q_{i}$ and corresponding final answers $a_{i}$. We also use Chain-of-Thought prompting Wei et al. (2022) with in-context examples in this process (details in Appendix A). In the second step, we sample multiple solutions for each $q_{i}$ using non-zero sampling temperatures, resulting in a synthetic dataset $\mathcal{D}_{S}=\{(q_{i},\{(\hat{r}_{ij},\hat{a}_{ij})\}_{j=1}^{k})\}$, where $k$ represents the number of sampled solutions. Here, $\hat{r}_{ij}$ denotes the $j$-th reasoning path (i.e., rationale) generated by the model for $q_{i}$, and $\hat{a}_{ij}$ is the model’s corresponding final answer. Incorrect solutions are then filtered out by comparing the sampled answers $\hat{a}_{ij}$ with the ground-truth answers $a_{i}$. Finally, we fine-tune the model on the filtered dataset $\tilde{\mathcal{D}}_{G}$ using supervised fine-tuning (SFT) to maximize the likelihood of generating reasoning paths $r$, optimizing the following objective:
$\mathbb{E}_{(q,r,a)\sim\tilde{\mathcal{D}}_{G}}\left[\log p_{\theta}(r,a|q)\right].$ (1)
Loss of Generalized Reasoning Ability during Self-Improvement
After applying the self-improvement framework to LLMs of various scales on in-domain (ID) datasets, we evaluate their performance on out-of-domain (OOD) reasoning datasets. The results, presented in Figure 2, reveal that while self-improvement enhances reasoning performance on ID datasets, it leads to a noticeable decline in performance on OOD datasets. This phenomenon suggests that although self-improvement improves metrics on ID reasoning tasks, it fails to enhance generalized reasoning capabilities and may even degrade them. We refer to this behavior as the emergence of Superficial Self-Improved Reasoners.
3.2 Investigating the Causes of Superficial Self-Improved Reasoners
Figure 3: The Layer Importance Scores of strong reasoning model Qwen2.5-1.5B-Math on BookCorpus (left) and MATH datasets (right). The middle layers are less important while the early and late layers are more important for reasoning (MATH). For non-reasoning task (BookCorpus) middle layers are more important.
Figure 4: The weight change for SFT Qwen2.5-1.5B with self-improvement MATH data (left) and fully post-training Qwen2.5-1.5B to Qwen2.5-1.5B-Math using real data with 700B tokens (right).
While numerous studies on catastrophic forgetting focus on analyzing and addressing OOD performance degradation in continual learning for learning simpler tasks, our work specifically targets the more challenging domain of mathematical reasoning in LLMs, with an emphasis on understanding the phenomenon of Superficial Self-Improved Reasoners. In this section, we identify the most critical layers for reasoning, analyze how their weights evolve during the self-improvement process, and provide an explanation for the emergence of Superficial Self-Improved Reasoners.
Layer Importance for Reasoning
To identify the most important weights in LLMs for reasoning, our objective is to determine and remove the weights that have the greatest impact on the model’s prediction, which can be measured by the resulting change in loss. We denote the linear weight matrix as $\mathbf{W}^{k,n}=\begin{bmatrix}W^{k,n}_{i,j}\end{bmatrix},$ where $k$ represents the modules (e.g., a key projection in the multi-head attention (MHA) or an up-projection in the feed-forward network (FFN)) within the $n$-th LLM layer. We quantify the importance of each weight by measuring the error introduced when the corresponding parameter is removed. Given an in-domain reasoning dataset $\mathcal{D}$, the importance score $I^{k,n}_{i,j}$ for the weight $W^{k,n}_{i,j}$ is defined as:
$I^{k,n}_{i,j}$ $=\left|\Delta\mathcal{L}(\mathcal{D})\right|$ (2)
$=\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W_{i,j}^{k,n}}W_{i,j}^{k,n}-\frac{1}{2}W_{i,j}^{k,n}H_{kk}W_{i,j}^{k,n}\right.$
$\quad\left.+\mathcal{O}\left(\|W_{i,j}^{k,n}\|^{3}\right)\right|.$
However, due to the significant computational cost associated with the large number of parameters in LLMs, we approximate the Hessian matrix $H_{kk}$ using the Fisher information matrix, following the approach in Ma et al. (2023). This allows us to approximate the second-order term $\frac{1}{2}W_{i,j}^{k,n}H_{kk}W_{i,j}^{k,n}$ as $\frac{1}{2}\sum_{j=1}^{N}\left(\frac{\partial\mathcal{L}(\mathcal{D}_{j})}{\partial W_{i}^{k}}W_{i}^{k}\right)^{2}$. By omitting the second-order derivative, the importance score $I^{k,n}_{i,j}$ is simplified to: $I^{k,n}_{i,j}\approx\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W^{k,n}_{i,j}}W^{k,n}_{i,j}\right|$. To assess the contribution of each layer to reasoning, we define the layer importance score as:
$I^{n}=\sum_{W^{k,n}_{i,j}}\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W^{k,n}_{i,j}}W^{k,n}_{i,j}\right|.$ (3)
We leverage this layer importance score $I^{n}$ to identify which layers contribute most significantly to reasoning tasks. As illustrated in Figure 3, the middle layers are less important while the early and late layers are more important for the reasoning (MATH) task. However, for the non-reasoning dataset BookCorpus, the middle layers are more important. This observation highlights the early and late layers as reasoning-critical layers, distinguishing their specialized function in reasoning.
Layer Weight Change after Self-Improvement
After fine-tuning the LLMs on reasoning data, the weights are updated, enabling the model to learn reasoning capabilities. We now analyze these weight changes. Let $\Delta\mathbf{W}^{n}$ represent the total weight change at the $n$-th layer after supervised fine-tuning (SFT), defined as:
$\Delta\mathbf{W}^{n}=\sum_{k}\left\|\mathbf{W}^{k,n}-\mathbf{W}^{k,n}_{\text{SFT}}\right\|,$ (4)
where $\mathbf{W}^{k,n}$ denotes the original $k$-th weight matrix and $\mathbf{W}^{k,n}_{\text{SFT}}$ is the fine-tuned weight matrix. Figure 4 illustrates the weight change $\Delta\mathbf{W}^{n}$ across different layers. For the self-improved model, the largest weight change occurs in the middle layers. In contrast, for the math model which is fully post-trained with stronger generalized reasoning capability, the most significant weight changes are concentrated in the early and late layers.
Model
Reasoning-
Critical Layer
Most Weight
Change Layer
Generalized
Reasoning Capability
Self-Improved Early, late Middle
Fully Post-trained Early, late Early, late $\checkmark$
Table 1: Comparison of self-improved model and fully post-trained math model.
Takeaway
By analyzing Figure 3 and Figure 4 (left), we observe that the middle layers (reasoning-trivial layers) are the least important for the strong reasoning capabilities of LLMs, yet these layers undergo the most significant updates during the self-improvement process. This phenomenon highlights a contradiction in how reasoning ability is acquired. If the model were solely learning generalized reasoning, the most substantial weight updates would occur in the early and late layers (reasoning-critical layers), as observed in fully post-trained math models with strong generalized reasoning capabilities, rather than in the middle layers.
This observation suggests that during self-improvement, the model does not exclusively enhance its reasoning ability but also exhibits a tendency to overfit the training data, effectively "memorizing" it. This overfitting behavior explains the improved performance on ID datasets while compromising the model’s generalization to OOD tasks. The performance comparison in Figure 2 further supports this conclusion. We summarize all experimental findings in Table 1, which leads to the following key insights: (i) during self-improvement on reasoning tasks, LLMs may show improved reasoning performance on ID tasks but lose generalized reasoning ability on OOD tasks; (ii) This phenomenon arises from a mismatch between the reasoning-critical layers and the layers with significant weight changes, suggesting that the model memorizes the training data rather than truly learning generalized reasoning capability.
4 Superficial Self-improved Reasoners Benefit from Iterative Model Merging
Figure 5: The overall framework. (a) The model generates chain-of-thought (CoT) answers for the given questions, and incorrect answers are filtered out using the ground-truth. The remaining correct answers are used for supervised fine-tuning (SFT) to self-improve the model. (b) IMM iteratively SFT the model and merges the self-improved models with the base model to balance reasoning enhancement and generalization.
Iterative Model Merging (IMM)
In this section, we propose Iterative Model Merging (IMM) to mitigate the Superficial Self-Improved Reasoners phenomenon, as illustrated in Figure 5. In the first self-improvement iteration, we self improve the original base model and merge the resulting SFT model $\boldsymbol{\theta_{SFT}^{0}}$ with the base model $\boldsymbol{\theta}$ to obtain the merged model $\boldsymbol{\theta_{m}^{0}}$. In each subsequent iteration $t$ ($t>0$), we continue the self-improvement process by fine-tuning the previously merged model $\boldsymbol{\theta_{m}^{t-1}}$. The resulting self-improved model $\boldsymbol{\theta_{SFT}^{t}}$ is then merged with the original base model to obtain the updated merged model $\boldsymbol{\theta_{m}^{t}}$. To formally describe this process, we define the parameter change $\boldsymbol{\delta^{t}}$ during each SFT iteration as follows:
$\boldsymbol{\delta^{t}}=\begin{cases}\boldsymbol{\theta^{t}_{SFT}}-\boldsymbol{\theta_{m}^{t}},&\text{if }t>0\text{, SFT merged model},\\ \boldsymbol{\theta^{t}_{SFT}}-\boldsymbol{\theta},&\text{if }t=0\text{, SFT base model}.\end{cases}$ (5)
We then incorporate DARE Yu et al. (2024a) to further process $\boldsymbol{\delta^{t}}$. DARE identifies parameter redundancy in LLMs, randomly masking parameter changes at a drop rate $p$ while scaling the remaining updates to improve the performance of the merged model. Denoting $\mathbf{m}\sim\text{Bernoulli}(p)$, DARE can be expressed as:
$\tilde{\boldsymbol{\delta^{t}}}=(1-\mathbf{m})\odot\boldsymbol{\delta^{t}},\quad\hat{\boldsymbol{\delta^{t}}}=\tilde{\boldsymbol{\delta^{t}}}/(1-p).$
By incorporating DARE into our iterative model merging framework, the final update for each iteration $t$ is given by:
$\boldsymbol{\theta_{m}^{t+1}}=\alpha\boldsymbol{\theta}+(1-\alpha)(\boldsymbol{\theta^{t}}+\boldsymbol{\hat{\delta^{t}}}),$ (6)
where $\alpha$ is a scaling parameter that controls the balance between the base model weights and the self-improved model weights.
Insights for IMM
The rationale behind model merging for generalized reasoning capability can be understood from two perspectives: (i) Based on the experimental observations in Section 3, the weights of reasoning-critical layers undergo significant changes during self-improvement, indicating that these layers are likely memorizing the training data. Given the blurred boundary between reasoning-critical and reasoning-trivial layers, it is plausible that middle layers also contribute to memorization, while late layers are partially involved in reasoning. As a result, excessive weight updates across all layers can lead to overfitting, especially when the training data is synthesized by the model itself. Model merging mitigates this overfitting by limiting weight changes. (ii) The base model retains strong generalization capabilities, while the self-improved model exhibits self-improved reasoning performance. Model merging combines the strengths of both, integrating the generalization ability of the base model with the reasoning improvements from the self-improved model.
5 Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of our proposed method. Specifically, our experiments aim to address the following research questions: (i) Can our method prevent model collapse on complex reasoning tasks during iterative self-improvement? (ii) How well does our method perform on OOD reasoning tasks? (iii) Can our method be extended from self-improvement to knowledge distillation from a stronger model?
Figure 6: The model performances on in-domain (ID) datasets performance. SFT n and Merge n denote the SFT model and merged model in the n-th iteration cycle. The model collapse happens from the first or second iteration for baselines while our method avoids it and achieves the best performance after model merging.
Figure 7: The model performances on out-of-domain (OOD) datasets. SFT n and Merge n denote the SFT model and merged model in the n-th iteration cycle. Baselines’ performances decrease on most datasets, while IMM can generally maintain the OOD performance compared with the original base model.
Figure 8: ID performance with different k for scaling up test-time-computing Pass@k.
5.1 Setup
Datasets
We train the model on MATH Hendrycks et al. (2021) and GSM-8K Cobbe et al. (2021) datasets correspondingly to evaluate the in domain reasoning ability of the model, while evaluate it on MAWPS Koncel-Kedziorski et al. (2016), SAT-Math Zhong et al. (2024) datasets to evaluate the out-of-domain reasoning ability.
Models
We include three LLMs at different scales (Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct Yang and et al. (2024) and Llama2-7B Touvron et al. (2023)) for self-improvement training. For the distillation experiment we include stronger teacher models Qwen2.5-7B-Instruct for distillation.
Baselines
We evaluate our method by comparing it with four baselines. First, we consider Vanilla (STaR Zelikman et al. (2022)) , which iteratively generates reasoning data following the procedure in Section 3 for self-improvement. Second, we include Data Mixture Shumailov et al. (2023), which mitigates performance degradation by mixing a portion of data from previous iterations. Third, we compare with Data Accumulation Gerstgrasser et al. (2024), which demonstrates that accumulating synthetic data across iterations can prevent model collapse.
Evaluation
[…]
We also provide the training and implementation details in Appendix B.
5.2 Main Results for Self-improvement
To answer the first research question, we conducted extensive experiments in a model collapse setting (iterative self-improvement) using two mathematical reasoning datasets, GSM8K and MATH. The results, shown in Figure 6, highlight that across three self-improvement iterations with three different LLMs, model collapse occurs in the first or second iteration for the baseline methods. In contrast, our method successfully avoids model collapse and achieves the best performance after applying model merging. Not only does our method significantly delay model collapse, but it also maintains superior performance across all iterations. Moreover, we observe that LLMs of all scales benefit from our model merging strategy, with smaller models suffering more severely from model collapse in the absence of this approach. Given the rising importance of test-time computing Snell et al. (2024), we further evaluate our method by generating multiple answers and measuring pass@$k$ accuracy. As shown in Figure 8, our method consistently improves performance as k increases and outperforms both the base models and the SFT models.
5.3 OOD Generalization Results
To answer the second research question, We evaluate the checkpoints from Section 5.2 using two out-of-domain (OOD) math reasoning datasets: SAT Math and MAWPS. The results, presented in Figure 7, show that while all other baselines suffer significant OOD performance degradation after iterative self-improvement, our method consistently restores performance after each model merging step and, in some cases, even surpasses the original base model. The only exception is the Qwen2.5-0.5B-Instruct model on the MAWPS dataset. We hypothesize that this dataset closely resembles the in-domain data, where extensive ID training significantly improves performance (from 12.8% to 62.9%), which may cause a slight degradation during the model merging process. Overall, these results demonstrate the great potential of our method, as it successfully mitigates the generalization drop commonly observed during SFT on specific tasks or datasets. Additional OOD datasets results can be found in Appendix F.
5.4 Distillation from Stronger Models
Considering self-improvement may only one of paradigm for LLM distillation, we extend our method to a broader field to answer the third research question. We distill stronger Qwen2.5-7B-Instruction model into the weaker Qwen2.5-1.5B-Instruction and Llama-2-7B models. The results in Table 2 demonstrate that the merged models consistently improve or maintain comparable performance on ID tasks, while often achieving significant improvements in OOD performance. This indicates that our method not only preserves task-specific performance but also enhances the model’s generalized reasoning ability when distilling from a stronger teacher model.
Student
Domain Datasets Base
SFT
Merged
[…]
6 Conclusion
This study identifies that self-improved LLM reasoners still lack generalized reasoning capability on OOD datasets, although they have better ID reasoning performances. We refer to this phenomenon as Superficial Self-Improved Reasoners and investigate its underlying causes. Our analysis reveals that during the self-improvement process, the layer weights change order doesn’t match the order of importance layers. This mismatch suggests that instead of solely learning to reason, the model also memorizes the training data. To address this issue, we propose Iterative Model Merging and extensive experiments demonstrate the effectiveness of our method: it not only mitigates model collapse but also have a generalized reasoning capability.
Limitations
The proposed Iterative Model Merging (IMM) method currently employs a fixed-weight merging mechanism between the original and self-improved models. However, more advanced strategies, such as dynamic or layer-adaptive merging, could provide further improvements. Additionally, although IMM has proven to effectively maintain generalized reasoning capabilities, it doesn’t investigate the strategy of mixing real and synthetic data appropriately, which could further enhance the trade-offs between reasoning improvement and generalization. We leave the exploration of advanced merging mechanisms and the optimal mixture ratio of real and synthetic data for future work.
Appendix A Chain of Thought Prompting for Data Synthesize
We use chain-of-thought prompting Wei et al. (2022) to generate answers. For MATH and GSM8K datasets, we both give 10 examples in the instructions for in-context generation. The prompting examples are given in Table 4.
Appendix B Training and implementation Details
[…]
Model Datasets Base SFT1 Merge1 SFT2 Merge2 SFT3 Merge3
Qwen2.5-0.5B-I SVAMP 7.3 35.8 5.9 21.6 1.3 40.1 9.8
ASDiv 8.7 51.4 3.7 30.7 2.8 46.7 17.6
MathQA 37.9 29.5 38.2 25.7 35.4 19.9 33.8
MMLU_stem 34.2 34.6 38.4 34.3 37.1 27.9 36.1
Qwen2.5-1.5B-I svamp 77.7 59 69.2 58.6 58.2 60.2 64.7
asdiv 82.8 72.5 76.4 64.8 59.6 70.8 73.4
MathQA 62.5 24.9 57.3 33.4 54.1 12.8 53.4
MMLU_stem 53.6 40.1 52.6 47.9 53.4 41.7 54.5
Llama2-7B svamp 39.6 30.1 38.0 35.1 39.0 33.5 38.5
asdiv 51.9 42.9 51.2 46.7 52.3 41.4 52.7
Table 3: OOD performance on additional reasoning datasets.
Appendix C Superficial Reasoning Finetuning Exists When Real Data is Limited
We also find that even using real but limited data, Superficial Reasoning Synthetic Finetuning still exists. As Figure 9 shows, the middle layers change most compared with the early and late layers, while Figure 3 already shows early and late layers are more important for reasoning. However, utilizing real data prevents the model from overfitting itself by using self-generated data. This is also verified by Figure 9: model’s reasoning layer (early and late layers) changed more (learn more reasoning capability) when training with real data, reasoning-trivial layers (middle layers)’s weight change is close to middle layers when training with synthetic data.
Figure 9: The weight change over layers for (i) Fintuning Qwen2.5-1.5B with synthetic MATH Hendrycks et al. (2021) dataset data and limited training data (7.5k real MATH training data)
Appendix D A bitter Lesson: Not All LLMs Can Self-improve
During our experiment, we also find that not all the LLMs can self-improve on reasoning tasks. If LLM’s performance decreases after SFT, then our method may not let the merged model have a better performance compared with the original model and the model after SFT. This usually happens when the original model already has a good performance (reasoning ability), and learned reasoning ability can’t offset the generalization loss. Unlike previous works that improve generalization and robustness via random smoothing Cohen et al. (2019); Yuan et al. (2024c) or regularization Virmaux and Scaman (2018); Ye et al. (2023); Yuan et al. (2024b), we utilize model merge to improve generalization and mitigate superficial self-improved reasoners phenomenon.
Appendix E The Connection to Catastrophic Forgetting
Figure 10: The layer importance score for Qwen2.5-1.5B base model on reasoning dataset MATH.
Catastrophic forgetting is a related but distinct phenomenon compared to superficial self-improved reasoners. Specifically, catastrophic forgetting refers to the loss of previously acquired knowledge when deep learning models are trained on new data. This issue occurs because model parameters are optimized based on the most recent training data, causing earlier learned representations to be dramatically overwritten.
While both catastrophic forgetting and superficial self-improved reasoners result in degraded performance due to further fine-tuning, their effects differ. After fine-tuning on new data, catastrophic forgetting results in a performance loss on previously learned tasks, whereas superficial self-improved reasoners result in diminished generalization capabilities on out-of-domain (OOD) tasks. This discrepancy arises because in catastrophic forgetting, fine-tuning on data for new tasks causes the model to lose knowledge from previous tasks. In contrast, superficial self-improved reasoners do not lead to forgetting too much past information but instead shift towards overfitting due to potentially biased knowledge, which may self-enhance along with the iteration of synthesizing new data and fine-tuning on it.
Appendix F Additional Experiments
Here we provide the additional experiment results for evaluating the layer importance for Qwen2.5-1.5B base model on reasoning datasets MATH. Similar to stronger reasoning model Qwen2.5-1.5B-Math, the importance layer for reasoning is early and late layers, as demonstrated in Figure 10. We also provide OOD performance on additional datasets SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), MathQA Amini et al. (2019) and MMLU-stem Hendrycks et al. (2020). IMM keeps the OOD reasoning capability as shown in Table 3.
Prompt for Generating GSM8K Answers
Below is an instruction that describes a task.
Write a response that appropriately completes the request like given examples below:
Question: Angelo and Melanie want to plan how many hours over the next week they should study together for their test next week.
They have 2 chapters of their textbook to study and 4 worksheets to memorize. They figure out that they should dedicate 3 hours to
each chapter of their textbook and 1.5 hours for each worksheet. If they plan to study no more than 4 hours each day, how many days
should they plan to study total over the next week if they take a 10-minute break every hour, include 3 10-minute snack breaks each
day, and 30 minutes for lunch each day?
A: Let’s think step by step.
Angelo and Melanie think they should dedicate 3 hours to each of the 2 chapters, 3 hours x 2 chapters = 6 hours total.
For the worksheets they plan to dedicate 1.5 hours for each worksheet, 1.5 hours x 4 worksheets = 6 hours total.
Angelo and Melanie need to start with planning 12 hours to study, at 4 hours a day, 12 / 4 = 3 days.
However, they need to include time for breaks and lunch. Every hour they want to include a 10-minute break, so 12 total hours x 10
minutes = 120 extra minutes for breaks.
They also want to include 3 10-minute snack breaks, 3 x 10 minutes = 30 minutes.
And they want to include 30 minutes for lunch each day, so 120 minutes for breaks + 30 minutes for snack breaks + 30 minutes for l
unch = 180 minutes, or 180 / 60 minutes per hour = 3 extra hours.
So Angelo and Melanie want to plan 12 hours to study + 3 hours of breaks = 15 hours total.
They want to study no more than 4 hours each day, 15 hours / 4 hours each day = 3.75
They will need to plan to study 4 days to allow for all the time they need.
The answer is 4
Question: Mark’s basketball team scores 25 2 pointers, 8 3 pointers and 10 free throws. Their opponents score double the 2 pointers
but half the 3 pointers and free throws. What’s the total number of points scored by both teams added together?
A: Let’s think step by step.
Mark’s team scores 25 2 pointers, meaning they scored 25*2= 50 points in 2 pointers.
His team also scores 6 3 pointers, meaning they scored 8*3= 24 points in 3 pointers
They scored 10 free throws, and free throws count as one point so they scored 10*1=10 points in free throws.
All together his team scored 50+24+10= 84 points
Mark’s opponents scored double his team’s number of 2 pointers, meaning they scored 50*2=100 points in 2 pointers.
His opponents scored half his team’s number of 3 pointers, meaning they scored 24/2= 12 points in 3 pointers.
They also scored half Mark’s team’s points in free throws, meaning they scored 10/2=5 points in free throws.
All together Mark’s opponents scored 100+12+5=117 points
The total score for the game is both team’s scores added together, so it is 84+117=201 points
The answer is 201
Question: Bella has two times as many marbles as frisbees. She also has 20 more frisbees than deck cards. If she buys 2/5 times
more of each item, what would be the total number of the items she will have if she currently has 60 marbles?
A: Let’s think step by step.
When Bella buys 2/5 times more marbles, she’ll have increased the number of marbles by 2/5*60 = 24
The total number of marbles she’ll have is 60+24 = 84
If Bella currently has 60 marbles, and she has two times as many marbles as frisbees, she has 60/2 = 30 frisbees.
If Bella buys 2/5 times more frisbees, she’ll have 2/5*30 = 12 more frisbees.
The total number of frisbees she’ll have will increase to 30+12 = 42
Bella also has 20 more frisbees than deck cards, meaning she has 30-20 = 10 deck cards
If she buys 2/5 times more deck cards, she’ll have 2/5*10 = 4 more deck cards.
The total number of deck cards she’ll have is 10+4 = 14
Together, Bella will have a total of 14+42+84 = 140 items
The answer is 140
Other 5 examples here …
### Instruction:
Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell
altogether in April and May?
### Response: Let’s think step by step.
Table 4: The CoT prompting examples for generate training data.